The radioisotope sodium-24, half-life 15 h, is used to measure the flow rate of salt water by irradiation of stable ?? with neutrons, suppose that we produce 5 micrograms of the isotope.

How long after a sample is placed in a reactor is it before the sample activity reaches 75% of the maximum activity? Assume the production of a single reactive species at a constant rate R sec-1, the decay rate λ, and there is no initial activity in the sample.
Answer #1

We use the equation:

A = A0 * 2-t/t0.5 , here A is the activity at time t ( 75% of A), A is the initial activity (A0), t0.5 is the half life

Using the values: 0.75 * A0 = A0 2-t/15

0.75 = 2-t/15

Solving, t = 6.22 Hours